OpenStudy (anonymous):

3 the square root of 4 times 3 the square root of 14

OpenStudy (anonymous):

i am going to make a guess that this is \[\sqrt[3]{4}\times \sqrt[3]{17}\] am i right?

OpenStudy (anonymous):

yes except it's 14 instead of 17 :)

OpenStudy (anonymous):

I have no idea what I'm doing do if someone could help that would be great!